# Decoherent histories with(out) objectivity in a (broken) apparatus

### v2

@@ Abstract
We characterize monitored quantum dynamics in a solvable model exhibiting a phase transition between a measurement apparatus and a scrambler. We show that approximate decoherent histories emerge in both phases with respect to a coarse-grained extensive observable. However, the apparatus phase, where quantum Darwinism emerges, is distinguished by the non-ergodicity of the histories and their correlation with the measured qubit, which selects an ensemble of preferred pointer states. Our results demonstrate a clear distinction between two notion of classicality, decoherent histories and environment-induced decoherence.
@@
The decoherence program [1, 2, 3, 4] is fundamentally motivated by a measurement problem, namely how to determine “what is being measured” when describing measurement as a unitary interaction between the system and the apparatus [5] . The key idea is that, the apparatus, and more generally, the environment) has many degrees of freedom, which allow to select a set of preferred “pointer states”, through the proliferation of classical records [6, 7]. In decoherence, and its refinement quantum Darwinism [8, 9, 10, 11, 12, 13, 14], classicality is defined as objectivity [15, 16, 17, 18], the redundant accessibility of records.
@@
Despite the intrinsic nature of classicality à la decoherent histories, many works have investigated decoherent histories in open quantum systems, motivated by the relation with the decoherence program [25, 26]. Some authors identify decoherent histories in the system (with the environment traced out) in the Markovian limit [27, 28], and make connection with quantum trajectories [29, 30, 31, 32]. Others treat the system and the environment as a whole, and derive decoherent histories (or a related branching structure) from objectivity [33, 34, 35]. This extensive literature fueled a folklore belief that decoherent histories require objectivity. This belief was challenged by recent works [36, 37, 38] showing evidence for (approximate) decoherent histories with respect to a coarse-grained and slowly varying observable, in an isolated macroscopic systems under chaotic evolution. Note that chaotic dynamics is known to scramble information, making them inaccessible for all practical purposes [39, 40, 41], which is the opposite of objectivity [42, 43]. Hence it seems that decoherent histories can exist independently of objectivity [44, 45, 46]. However the mechanisms appear entirely different, and the relation between the two notions of classicality awaits clarification.
In this Letter, we characterize decoherent histories in a solvable model [47, 48] where objectivity can be turned on or off: the model has a transition between a functioning apparatus and a scrambler (Fig. 1). We will show that decoherent histories emerge in both phases with respect to a coarse-grained observable, the coarse-graining being the essential common mechanism (Fig. 2). Nevertheless, we shall distinguish the phases by the (non)-ergodicity of the histories, and their correlation with the measured qubit, described in terms of a “pointer states ensemble” that we shall introduce (Fig. 3).
@@
This is a situation where the aforementioned measurement basis problem is the most acute: it is impossible to know in which basis the measurement took place [6, 7].
In our apparatus model, $|\Psi\rangle$ is the initial state of a cascade process involving more and more qubits, bridging micro- and macroscopic realms [49, 50, 51]. Starting with the qubit $A$, at each time step, every existing qubit interacts with a new qubit (in a pure state), so that there are $N_{t}=2^{t}$ qubits after $t$ steps. During all this, $S$ remains intact. Such an inflationary dynamics is known as “concatenation” in the error correction code literature [52, 53, 54], and often represented diagrammatically by a tree, see Fig. 1 [55, 56]. Every node corresponds to an isometry that outputs two qubits from one, $v:\mathbb{C}^{2}\to\mathbb{C}^{2}\otimes\mathbb{C}^{2}$, such that $v^{\dagger}v=I$. The evolution from time step $t$ to $t+1$ is given by $V_{t+1,t}=v^{\otimes 2^{t}}$. We also denote $V_{t,s}:=V_{t,t-1}V_{t-1,t-2}\dots V_{s+1,s}$ for $t\geq s$. We will focus on the following concrete family of models:
@@
where $N_{t}=2^{t}$ and $Z_{j}$ acts on the $j$-th qubit at time $t$. At $\theta=0$, $\mathcal{Q}_{t}$ is also perfectly correlated with $Z_{S}$, the $Z$ operator on $S$. Exploiting the tree structure (see below and [57] for methods), we calculate the correlation between $O^{\prime}_{S}$ (for any $O^{\prime}$) and $\mathcal{Q}_{t}$ for any $\theta$. As $t\to\infty$, we find the following scaling behavior
@@
We remark that our expanding tree model is a special case of the multi-scale entanglement renormalization Ansatz (MERA) tensor networks [59, 60, 61, 62], which are known to generate scale invariant states. In that context, the exponent $x_{1}$ above is the lightest scaling dimension. Interestingly, the threshold phenomena at $x_{1}=1/2$ appeared in various recently studied scale invariant “inference” problems [63, 64, 65].
@@
In the sense of (12), classicality à la decoherent histories amounts to non-disturbance by third-party measurements (which is also part of macro-realism [66]). Namely, classicality emerges if the existence of measurements in $t\notin\mathbf{t}$ do not affect the outcome distribution of the ones in $t\in\mathbf{t}$. This is a reasonable way to define “classical” since only classical systems admit ideal measurements that reveal their state without changing it.
@@
Figure 2: Decoherent histories probe $\|\Delta^{\mathbf{t}}\|_{1}$ with $\mathbf{t}=\{L+\tau\}$, as function of $\tau$ (measurement starting time), for (a) $\theta=0.15\pi<\theta_{c}$ and (b) $\theta=0.3\pi>\theta_{c}$. Approximate decoherent histories emerge for coarse-grained measurements ($\Gamma=0.1$) and $\tau\gg 1$, but not for $\tau\sim 1$ or for fine-grained measurements (fg, $\Gamma\eta_{t}=0.1$). Insets (i, ii) show the pdf $p^{\mathbf{t}}$ (no third-party) and marginal distributions in non-decoherent scenarios ($T=11$). Straight lines show decay rate of $\|\Delta^{\mathbf{t}}\|_{1}$ (14), and (55) in [57]. (a) For $\theta<\theta_{c}$, $\|\Delta^{\mathbf{t}}\|_{1}\sim\Gamma^{-2}2^{2\tau(x-1)}$ with no $L$ scaling. (b) For $\theta>\theta_{c}$, : $\|\Delta^{\mathbf{t}}\|_{1}\sim\Gamma^{-2}2^{-T}\max(1,2\cos\theta)^{2L}$. Inset (iii): $L$-dependence is shown for two $\theta>\theta_{c}$ ($\tau=6$).
@@
The emergence of decoherent histories crucially relies on coarse-graining, that is, measuring an extensive quantity with $\Gamma=O(1)$ relative precision, see (13). Had we let $\Gamma\sim O(1/\eta_{t})$, $\|\Delta^{\mathbf{t}}\|_{1}\not\to 0$ in any limit (see blue crosses in Fig. 2). Such a measurement with $O(1)$ absolute precision would disturb the dynamics even in a macroscopic system, revealing its underlying quantum nature. Ref. [70] also pointed out the importance of measurement imprecision for classicality in macroscopic systems, in terms of the Leggett-Garg inequality [66]. Interestingly, in our model, this inequality can be violated at arbitrarily late time in both phases, by operators of type $Q_{t}=(aX_{j}+bY_{j}+cZ_{j})^{\otimes 2^{t}},Q_{t}^{2}=\mathbf{1}$ [57]. Again, such an observable reveals the many-body coherence of the model inaccessible to coarse-grained ones.
@@
where $\xi_{t}$ are i.i.d standard Gaussian and $\mathcal{M}$ is a random variable with the same distribution as $\mathcal{Q}_{t}/\eta_{t}$ in the long time limit (this distribution is non-Gaussian, see Fig. 2):
@@
Pointer states ensemble. We come back to the question: what information on the measured qubit $S$ is inferred from an apparatus history? This can be described by the ensemble of the conditional state $\rho_{\vec{m}}$ of $S$ after a coarse-grained monitoring [using (13)] of the apparatus, weighted by Born’s rule $p_{m}$ [48, 71, 72, 73], which we call the pointer states ensemble. This ensemble is different from and “softer” than the usual definitions of pointer states [6, 28, 74]. In general, it is made of mixed states that do not form a basis. Yet, they constitute a “decomposition of unity” like a basis: $\sum_{\vec{m}}p_{\vec{m}}\rho_{\vec{m}}=I$ [48]. Moreover, it captures an imperfect form of apparatus super-selection, as we now illustrate.
@@
Conclusion. We distinguished two notions of classicality. Defined via decoherent histories, its emergence in a macroscopic object relies only on coarse-graining and is likely generic. Yet, classicality à la quantum Darwinism is about the particular macro-microscopic correlation as realized during a quantum measurement. An apparatus is a macroscopic object, but not every macroscopic object is an apparatus. This statement may sound obvious, yet is nontrivial to demonstrate: we only did it in a somehow unusual “inflationary” model. Indeed, it would be desirable to confront objectivity and decoherent histories in conventional lattice models beyond small sizes. There, stabilizing objectivity is nontrivial, and plausibly requires some form of symmetry breaking [42, 76, 43, 51]. Meanwhile, implication of our study for the issue of classicality in cosmology [77] awaits exploration.
@@ Acknowledgements.
We thank Serge Florens and Vincent Vennin for useful discussion and collaboration on related topics.

### v3

@@ Abstract
We characterize monitored quantum dynamics in a solvable model exhibiting a phase transition between a measurement apparatus and a scrambler. We show that approximate decoherent histories emerge in both phases with respect to a coarse-grained extensive observable. However, the apparatus phase, where quantum Darwinism emerges, is distinguished by the non-ergodicity of the histories and their correlation with the measured qubit, which selects an ensemble of preferred pointer states. Our results demonstrate a clear distinction between two notions of classicality, decoherent histories and environment-induced decoherence.
@@
The decoherence program [1, 2, 3, 4] is fundamentally motivated by a measurement problem, namely how to determine “what is being measured” when describing measurement as a unitary interaction between the system and the apparatus [5]. The key idea is that the apparatus (more generally, the environment) has many degrees of freedom, which allow to select a set of preferred “pointer states”, through the proliferation of classical records [6, 7]. In decoherence, and its refinement quantum Darwinism [8, 9, 10, 11, 12, 13, 14], classicality is defined as objectivity [15, 16, 17, 18], the redundant accessibility of records.
@@
Despite the intrinsic nature of classicality à la decoherent histories, many works have investigated decoherent histories in open quantum systems, motivated by the relation with the decoherence program [25, 26]. Some authors identify decoherent histories in the system (with the environment traced out) in the Markovian limit [27, 28], and make connection with quantum trajectories [29, 30, 31, 32]. Others treat the system and the environment as a whole, and derive decoherent histories (or a related branching structure) from objectivity [33, 34, 35]. This extensive literature fueled a folklore belief that decoherent histories require objectivity. This belief was challenged by recent works [36, 37, 38] showing evidence for (approximate) decoherent histories with respect to a coarse-grained and slowly varying observable, in an isolated macroscopic system under chaotic evolution. Note that chaotic dynamics is known to scramble information, making it inaccessible for all practical purposes [39, 40, 41], which is the opposite of objectivity [42, 43]. Hence it seems that decoherent histories can exist independently of objectivity [44, 45, 46]. However the mechanisms appear entirely different, and the relation between the two notions of classicality await clarification.
In this Letter, we characterize decoherent histories in a solvable model [47, 48] where objectivity can be turned on or off: the model has a transition between a functioning apparatus and a scrambler (Fig. 1). We will show that decoherent histories emerge in both phases with respect to a coarse-grained observable, the coarse-graining being the essential common mechanism (Fig. 2). Nevertheless, we shall distinguish the phases by the (non)-ergodicity of the histories, and their correlation with the measured qubit, described in terms of a “pointer state ensemble” that we shall introduce (Fig. 3).
@@
This is a situation where the aforementioned measurement basis problem is the most acute: it is impossible to know in which basis the measurement took place [6, 7]. This issue is to be fixed by the presence of an environment or a more complex apparatus model, as we define now.
In our model, $|\Psi\rangle$ is the initial state of a cascade process involving more and more qubits, bridging micro- and macroscopic realms [49, 50, 51]. Starting with the qubit $A$, at each time step, every existing qubit interacts with a new qubit (in a pure state), so that there are $N_{t}=2^{t}$ qubits after $t$ steps. During all this, $S$ remains intact. Such an inflationary dynamics is known as “concatenation” in the error correction code literature [52, 53], and often represented diagrammatically by a tree, see Fig. 1 [54, 55]. Every node corresponds to an isometry that outputs two qubits from one, $v:\mathbb{C}^{2}\to\mathbb{C}^{2}\otimes\mathbb{C}^{2}$, such that $v^{\dagger}v=I$. The evolution from time step $t$ to $t+1$ is given by $V_{t+1,t}=v^{\otimes 2^{t}}$. We also denote $V_{t,s}:=V_{t,t-1}V_{t-1,t-2}\dots V_{s+1,s}$ for $t\geq s$. We will focus on the following concrete family of models:
@@
where $N_{t}=2^{t}$ and $Z_{j}$ acts on the $j$-th qubit at time $t$. At $\theta=0$, $\mathcal{Q}_{t}$ is also perfectly correlated with $Z_{S}$, the $Z$ operator on $S$. Exploiting the tree structure (see below and [56] for methods), we calculate the correlation between $O^{\prime}_{S}$ (for any $O^{\prime}$) and $\mathcal{Q}_{t}$ for any $\theta$. We find the following asymptotic behavior (that is, the dependence on $t$ in the $t\to\infty$ limit):
@@
We remark that our expanding tree model is a special case of the multi-scale entanglement renormalization Ansatz tensor networks [58, 59, 60, 61], which are known to generate scale invariant states. In that context, the exponent $x_{1}$ above is the lightest scaling dimension. Interestingly, the threshold phenomena at $x_{1}=1/2$ appeared in various recently studied scale invariant “inference” problems [62, 63, 64].
@@
In the sense of (12), classicality à la decoherent histories amounts to non-disturbance by third-party measurements (which is also part of macro-realism [65]). Namely, classicality emerges if the existence of measurements in $t\notin\mathbf{t}$ does not affect the outcome distribution of the ones in $t\in\mathbf{t}$. This is a reasonable way to define “classical” since only classical systems admit ideal measurements that reveal their state without changing it.
@@
Figure 2: Decoherent histories probe $\|\Delta^{\mathbf{t}}\|_{1}$ with $\mathbf{t}=\{L+\tau\}$, as function of $\tau$ (measurement starting time), for (a) $\theta=0.15\pi<\theta_{c}$ and (b) $\theta=0.3\pi>\theta_{c}$. Approximate decoherent histories emerge for coarse-grained measurements ($\Gamma=0.1$) and $\tau\gg 1$, but not for $\tau\sim 1$ or for fine-grained measurements (fg, $\Gamma\eta_{t}=0.1$). Insets (i, ii) show the pdf $p^{\mathbf{t}}$ (no third-party) and marginal distributions in non-decoherent scenarios ($T=11$). Straight lines show decay rate of $\|\Delta^{\mathbf{t}}\|_{1}$ (14), and (55) in [56]. (a) For $\theta<\theta_{c}$, $\|\Delta^{\mathbf{t}}\|_{1}\sim\Gamma^{-2}2^{2\tau(x-1)}$ with no $L$ scaling. (b) For $\theta>\theta_{c}$, : $\|\Delta^{\mathbf{t}}\|_{1}\sim\Gamma^{-2}2^{-T}\max(1,2\cos\theta)^{2L}$. Inset (iii): $L$ dependence is shown for two $\theta>\theta_{c}$ ($\tau=6$).
@@
The emergence of decoherent histories crucially relies on coarse-graining, that is, measuring an extensive quantity with $\Gamma=O(1)$ relative precision, see (13). Had we let $\Gamma\sim O(1/\eta_{t})$, $\|\Delta^{\mathbf{t}}\|_{1}\not\to 0$ in any limit (see blue crosses in Fig. 2). Such a measurement with $O(1)$ absolute precision would disturb the dynamics even in a macroscopic system, revealing its underlying quantum nature. Ref. [69] also pointed out the importance of measurement imprecision for classicality in macroscopic systems, in terms of the Leggett-Garg inequality [65]. Interestingly, in our model, this inequality can be violated at arbitrarily late times in both phases, by operators of type $Q_{t}=(aX_{j}+bY_{j}+cZ_{j})^{\otimes 2^{t}},Q_{t}^{2}=\mathbf{1}$ [56]. Again, such an observable reveals the many-body coherence of the model inaccessible to coarse-grained ones.
@@
where $\xi_{t}$ are independent identically distributed standard Gaussian and $\mathcal{M}$ is a random variable with the same distribution as $\mathcal{Q}_{t}/\eta_{t}$ in the long time limit (this distribution is non-Gaussian, see Fig. 2):
@@
Pointer states ensemble. We come back to the question of what information on the measured qubit $S$ is inferred from an apparatus history. This can be described by the ensemble of the conditional state $\rho_{\vec{m}}$ of $S$ after a coarse-grained monitoring [using (13)] of the apparatus, weighted by Born’s rule $p_{m}$ [48, 70, 71, 72], which we call the pointer states ensemble. This ensemble is different from and “softer” than the usual definitions of pointer states [6, 28, 73]. In general, it is made of mixed states that do not form a basis. Yet, they constitute a “decomposition of unity” like a basis: $\sum_{\vec{m}}p_{\vec{m}}\rho_{\vec{m}}=I$ [48]. Moreover, it captures an imperfect form of apparatus super-selection, as we now illustrate.
@@
Conclusion. We distinguished two notions of classicality. Defined via decoherent histories, its emergence in a macroscopic object relies only on coarse-graining and is likely generic. Yet, classicality à la quantum Darwinism is about the particular macro-microscopic correlation as realized during a quantum measurement. An apparatus is a macroscopic object, but not every macroscopic object is an apparatus. This statement may sound obvious, yet is nontrivial to demonstrate: we only did it in a somehow unusual “inflationary” model. Indeed, it would be desirable to confront objectivity and decoherent histories in conventional lattice models beyond small sizes. There, stabilizing objectivity is nontrivial, and plausibly requires some form of symmetry breaking [42, 75, 43, 51]. Meanwhile, the implication of our study for the issue of classicality in cosmology [76] awaits exploration.
@@ Acknowledgements.
We thank Serge Florens and Vincent Vennin for useful discussion and collaboration on related topics. The data that support the findings of this article are openly available [77].